# TemporalBench: Benchmarking Fine-grained Temporal Understanding for Multimodal Video Models

### v1

@@ Appendix C Per subset Results for Short and Long Video QA under Binary Accuracy (BA)
Appendix D More Results with Extended Frames
Table 10: TemporalBench performance of various models under binary QA and multiple binary QA setting for short question answering with different number of frames.
Model Frames Per Video Multiple Binary Accuracy (%) Binary QA Accuracy (%)
Human - 67.9 89.7
Random Chance - 9.5 50.0
XCLIP 8 12.9 51.6
ImageBind 2 14.0 53.0
LanguageBind 8 14.5 52.8
GPT-4o 64 38.0 76.0
32 38.2 75.9
16 38.5 75.7
8 37.3 75.1
4 35.8 74.4
2 33.2 72.7
1 28.4 70.0
0 26.5 67.7
Gemini-1.5-Pro 1FPS 26.6 67.5
0 16.1 58.1
Claude-3.5-Sonnet 16 23.5 65.4
8 23.6 65.5
4 23.1 64.8
2 21.2 61.8
1 18.4 58.4
InternLM-XC25 1FPS 17.9 58.8
LLaVA-NeXT-Video-34B-DPO 32 22.0 64.0
16 21.8 63.7
8 21.4 63.3
4 20.7 63.0
2 19.9 61.8
1 18.8 60.5
LLaVA-NeXT-Video-7B-DPO 32 17.2 59.5
16 22.3 64.0
8 23.6 65.1
4 22.9 64.2
2 21.4 63.0
1 19.0 62.0
VideoLLaVA 8 25.5 67.1
Phi-3.5-Vision-Instruct 32 15.5 56.7
16 15.9 57.2
8 15.9 57.4
4 15.5 57.5
2 16.9 58.0
1 16.4 57.7
Qwen2-VL-7B 32 24.7 64.4
16 23.6 63.2
8 21.0 60.9
4 19.2 59.5
2 17.6 57.8
Qwen2-VL-72B 32 38.3 75.8
16 36.8 74.6
8 33.8 73.0
4 31.0 71.4
2 27.3 69.1
MiniCPM-V-2.6 1FPS 21.4 62.3
LLaVA-1.5-13B 1 13.1 55.7
LLaVA-1.5-7B 1 18.3 60.5
Phi-3-Vision 1 15.1 54.4
Yi34B 0 18.7 59.9
Vicuna7B-1.5 0 10.4 50.55
Flan-T5-XL 0 17.9 57.9
Flan-T5-XXL 0 15.1 55.1
Appendix E Data Annotation Platform
Positive Captions
Negative Captions
Figure 11: Positive caption refinement platform.
Figure 12: Negative caption annotation platform.

### v2

@@ Appendix C Per subset Results for Short and Long Video QA under Binary Accuracy (BA)
Appendix D More Results with Extended Frames
Table 8: TemporalBench performance of various models under binary QA accuracy (BA) and multiple binary QA accuracy (MBA) setting for short and long question answering with different number of frames. “Overall” denotes the average performance of short and long video QA performance.
Model # Frames Overall MBA Overall BA Short MBA Short BA Long MBA Long BA Captioning
Human Performance - - - 67.9 89.7 - - -
Random Chance - 9.5 50.0 9.5 50.0 9.5 50.0 -
XCLIP 8 12.0 51.7 12.9 51.6 11.1 51.7 -
ImageBind 2 12.4 52.0 14.0 53.0 10.7 51.0 -
LanguageBind 8 13.3 52.2 14.5 52.8 12.0 51.6 -
GPT-4o 64 35.4 73.3 38.0 76.0 32.7 70.5 63.5
32 32.9 71.5 38.3 75.9 27.4 67.0 63.2
16 34.3 72.8 38.5 75.7 30.1 69.8 61.3
8 32.9 72.0 37.4 75.1 28.3 68.8 60.3
4 31.9 71.2 35.8 74.4 28.0 68.0 58.8
2 30.3 70.3 33.3 72.7 27.3 67.8 55.3
1 26.5 67.4 28.4 70.0 24.5 64.7 52.3
0 27.4 67.7 26.5 67.7 28.2 67.6 -
Gemini-1.5-Pro 1FPS 25.7 66.4 26.6 67.5 24.7 65.2 56.5
0 18.7 60.2 16.1 58.1 21.2 62.2 -
Claude-3.5-Sonnet 16 23.2 64.2 23.5 65.9 22.9 62.4 54.1
8 24.1 65.1 23.6 65.5 24.5 64.6 53.1
4 23.2 64.2 23.1 64.8 23.3 63.6 51.9
2 21.1 62.2 21.2 61.9 20.9 62.4 48.2
1 18.7 58.9 18.4 58.5 18.9 59.3 41.0
Qwen2-VL-72B 32 31.7 70.2 38.3 75.8 25.0 64.5 56.1
16 31.5 70.1 36.9 74.6 26.1 65.5 54.1
8 30.1 68.9 34.0 73.1 26.2 64.7 51.4
4 28.6 68.4 31.2 71.5 26.0 65.3 48.3
2 27.3 67.6 27.5 69.2 27.1 66.0 43.9
Qwen2-VL-7B 32 21.8 62.1 24.7 64.4 18.8 59.7 51.9
16 21.2 61.5 23.6 63.3 18.7 59.7 50.3
8 19.2 59.4 21.1 61.1 17.2 57.7 48.4
4 17.4 58.5 19.3 59.5 15.4 57.5 46.1
2 16.4 56.9 17.7 57.8 15.0 56.0 42.0
LLaVA-OneVision-72B 32 26.6 66.6 30.7 70.5 22.4 62.7 53.9
16 27.2 67.3 32.1 71.2 22.3 63.4 54.2
8 28.1 67.9 33.0 72.1 23.1 63.6 55.0
4 27.6 67.3 31.4 71.2 23.8 63.4 54.2
2 25.7 66.3 29.2 69.6 22.1 63.0 51.1
1 23.3 64.1 27.1 67.9 19.5 60.2 48.6
LLaVA-OneVision-7B 32 18.7 59.4 21.2 61.9 16.2 56.9 50.1
16 17.9 58.8 20.1 60.9 15.6 56.6 50.4
8 17.3 57.8 19.5 59.9 15.0 55.7 50.2
4 16.4 56.3 18.9 58.9 13.9 53.7 49.7
2 14.8 54.4 16.8 56.1 12.7 52.7 47.1
1 12.0 51.4 13.4 53.3 10.6 49.5 44.1
LLaVA-NeXT-Video-34B 32 19.9 61.1 22.0 64.0 17.7 58.2 53.1
16 20.3 61.1 21.8 63.7 18.7 58.4 53.3
8 20.5 61.9 21.4 63.3 19.5 60.4 53.4
4 20.4 61.7 20.7 63.0 20.0 60.3 52.5
2 19.9 61.2 20.0 61.8 19.7 60.6 48.9
1 19.0 59.8 19.0 60.5 18.9 59.1 46.2
LLaVA-NeXT-Video-7B 32 15.9 57.1 17.3 59.5 14.5 54.7 51.6
16 19.3 59.9 22.4 64.0 16.1 55.7 49.9
8 20.5 61.2 23.6 65.1 17.3 57.2 50.1
4 20.0 60.7 23.0 64.2 17.0 57.2 49.2
2 19.2 60.3 21.5 63.1 16.8 57.4 46.8
1 17.6 59.1 19.1 62.0 16.1 56.1 44.0
InternLM-XC2.5 1FPS 16.8 57.3 17.9 58.8 15.6 55.8 52.4
VideoLLaVA 8 20.3 61.6 25.5 67.1 15.1 56.0 46.0
MiniCPM-V2.6 1FPS 20.4 61.3 21.4 62.3 19.3 60.3 47.2
Phi-3.5-Vision 32 14.1 54.3 15.6 56.8 12.6 51.7 48.4
16 14.7 55.1 15.9 57.2 13.5 53.0 48.9
8 14.9 55.6 15.9 57.4 13.8 53.7 48.3
4 15.0 56.0 15.5 57.5 14.5 54.5 44.0
2 15.5 56.2 16.9 58.0 14.1 54.4 42.9
1 14.5 55.9 16.5 57.7 12.5 54.0 42.1
MA-LMM 4 9.1 47.4 9.2 48.0 9.0 46.9 38.7
M3 6 13.3 54.7 14.8 56.4 11.8 53.1 47.8
LLaVA-1.5-13B 1 13.7 55.1 13.1 55.7 14.2 54.5 47.9
LLaVA-1.5-7B 1 15.3 56.8 18.3 60.5 12.3 53.2 45.7
LLaVA-NeXT-34B 1 19.0 60.5 18.0 60.5 19.9 60.5 49.1
Phi-3-Vision 1 15.4 55.2 15.1 54.4 15.6 56.0 42.0
Gemini-1.5-Pro 0 18.6 60.1 16.1 58.1 21.2 62.2 -
Yi-34B 0 18.5 59.7 18.7 59.9 18.4 59.5 -
Vicuna7b-1-5 0 10.1 50.8 10.4 50.5 9.9 51.1 -
Flan-T5-XL 0 18.6 59.0 17.9 57.9 19.4 60.1 -
Flan-T5-XXL 0 15.9 56.0 15.1 55.1 16.7 56.9 -
Appendix E Data Annotation Platform
Positive Captions
Negative Captions
Figure 11: Positive caption refinement platform.
Figure 12: Negative caption annotation platform.